A snowman is built with three spheres made of packed snow. The top sphere has half the diameter of the middle sphere. The middle sphere has half the diameter of the bottom sphere. The top sphere weighs 5 pounds. What is the total weight of the snowman?
a. 30 pounds
b. 65 pounds
c. 125 pounds
d. 250 pounds
e. 365 pounds

Answer by mananth(16949)


since the radius are in the ratio of 1:2:4
the volume will be 1:8:64
the top one is 5 pounds
the second will be 8*5 = 40 pounds
the bottom one will be 64*5 = 320 pounds
So total weight will be 5+40+320= 365 pounds.
..